OpenStudy (anonymous):

Please help.. The radioactive substance strontium-90 has a half-life of 28 years. In other words, it takes 28 years for half of a given quantity of strontium-90 to decay to a non-radioactive substance. The amount of radioactive strontium-90 still present after t years is modeled by the expression 80 * 2^-t/28 grams. Evaluate the expression for t = 0, t = 28 and t = 56. What does each value of the expression represent?

OpenStudy (anonymous):

at t=0 80 * 2^0/28 =80 it represent the incial amount at t=28 80 * 2^-28/28 =40 it represents the amount left after 28 years (half life) at t= 56 80 * 2^-56/28=20 it represents the amount left after 56 years

OpenStudy (anonymous):

thank you very much!

OpenStudy (anonymous):

yw